According to a​ website, the average cost of a stay in a hospital for a particular region was 12,500.

Assume that this cost is exponentially distributed.

a. Determine the probability that a randomly selected​ patient's stay in a hospital in this region will cost more than

​$13,000.

b. Calculate the probability that a randomly selected​ patient's stay in a hospital in this region will cost less than

$9,700.
Answer #1

Answer

(A) given that mean = 12500

we have to find P(X> 13000)

using the formula

P(X>x) = e^(-x/mean)

= e^(-13000/12500)

= 0.3535

(B)

given that mean = 12500

we have to find P(X<9700)

using the formula

P(X<x) = 1 - e^(-x/mean)

= 1 - e^(-9700/12500)

= 1-0.4602

=0.5398